### v2

@@ 6.2 Perspectives for future work
A future work will relax the definition of $\gamma(x)$ as a tractable global solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for all $x\in X$. We plan to alter the POf to allow for the access to only an approximation of a local solution. This would be representative of most practical cases, where all subproblems are solved numerically by an optimization method which only approximates a local solution. We conducted some tests following this idea on some problems from [9, Chapter 7] by defining, for all $x\in X$$\gamma(x)$ as the output of a numerical method solving Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). This approach seems promising and works well in our experiments, although the theory is complex so we are still working to simplify it. This future work may also determine if, given a point $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{ref}}$), the point $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{ini}}$). This would allow for an adaptation of Theorems 1 and 2 where Reformulation ($\mathbf{P}_{\texttt{ref}}$) is tackled with an algorithm that return points satisfying only necessary optimality conditions.
Given a collection of partition sets of $\mathbb{Y}$, there is a flexibility in the way to index them. This choice of indexation impacts the shape of $\Phi$. Then, we may fine-tune the indexation of the partition sets of any partition of $\mathbb{Y}$ so that the resulting Reformulation ($\mathbf{P}_{\texttt{ref}}$) is solved as easily as possible.
@@
Reformulation ($\mathbf{P}_{\texttt{ref}}$) is currently defined according to an extreme barrier [3], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible. We observed in [9, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value +$\infty$+ ∞. A progressive barrier [4] may be more efficient. We have initiated some research to associate an infeasibility metric to all points $x\in\mathbb{X}$ and use it in the progressive barrier. For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{ref}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). In all the problems we tested in [9, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is significantly easier with such a progressive barrier than than with an extreme barrier.
@@
This framework may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for any $x\in X$). We plan to develop more advanced hybrid methods. Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{ini}}$) directly by optimizing jointly the partition set to consider and the point to consider into the set.
Finally, the class of composite problems illustrated in Section 4 is not the only class where the POf might be considered. A noticeable example lies in partially separable problems, where fixing some key variables makes the subproblem fully separable. Another example consists in problems where each variable has little individual influence on the objective value but their joint influence is much greater, similarly to the state variables in optimal control theory. As the theory is now stated, applying the POf requires only to follow the content of Section 3. We plan to reword our former work [10] accordingly.

### v3

@@ 6.2 Perspectives for future work
Two lines of improvements appear for extending the POf. The first concerns the oracle function $\gamma$, as it is unrealistic to assume that a global solution to each subproblem is computable. A future work will define $\gamma$ as, for all $x\in X$, an approximation of a local solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$). This would be representative of most practical cases. We conducted some tests following this idea on some optimal control problems in [10, Chapter 7] by defining, for all $x\in X$, $\gamma(x)$ as the output of a well-suited numerical method solving Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) (chosen according to acknowledged guidelines from the dedicated literature). This approach seems promising and works well in our experiments, although the theory is technical so we are still working to simplify it. We may also determine if, given a point $x^{*}\in X$ satisfying necessary optimality conditions for Reformulation ($\mathbf{P}_{\texttt{r}}$), the point $\gamma(x^{*})$ satisfies necessary optimality conditions for Problem ($\mathbf{P}_{\texttt{i}}$). This would allow tackling Reformulation ($\mathbf{P}_{\texttt{r}}$) with any DFO algorithm that returns points satisfying only necessary optimality conditions, instead of imposing the covering step to ensure local optimality.
@@
Another possible improvement concerns the solving process of Reformulation ($\mathbf{P}_{\texttt{r}}$). Currently, Reformulation ($\mathbf{P}_{\texttt{r}}$) is defined according to an extreme barrier [4], since $\Phi(x)\triangleq+\infty$ for all $x\in\mathbb{X}$ such that Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is infeasible. We observed in [10, Chapter 7] that an extreme barrier may lead to an important number of evaluated points returning the value $\infty$. A progressive barrier [5] may be more efficient. We have initiated some research to associate an infeasibility metric to all points $x\in\mathbb{X}$ and use it in the progressive barrier. For all $x\in\mathbb{X}$, our infeasibility associated to $x$ with respect to Reformulation ($\mathbf{P}_{\texttt{r}}$) is the infimum over all $y\in\mathbb{Y}(x)$ of the infeasibility of $y$ with respect to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$). In all the problems we tested in [10, Chapter 7], solving Reformulation ($\mathbf{P}_{\texttt{r}}$) is significantly easier with such a progressive barrier than with an extreme barrier.
@@
Beyond these two main lines of improvement, we also plan to develop more advanced algorithms for the DFPOm. This may be seen as a prototypical hybrid method using either a DFO algorithm (to solve Reformulation ($\mathbf{P}_{\texttt{r}}$)) and other classes of algorithms (to solve Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) for any $x\in X$). Roughly speaking, we may solve Problem ($\mathbf{P}_{\texttt{i}}$) by optimizing jointly the partition set to consider and the point to consider into the set. We may also learn or dynamically adapt the partition and its indexation to improve the behaviour of $\Phi$ and the computational cost of $\gamma$.
Finally, Section 4 does not enclose all problems where the POf is useful. As the theory is now stated, applying the POf requires only to follow Section 3. We plan to reword our work [11] about discontinuous optimal control problems accordingly. Another direction is related to parametric optimization [18, 30], since our results are directly transferrable to this context. We also plan to tackle counterfactuals for contextual optimization from machine learning [28, 31, 33], which may relate well to the POf.